# Disturbance-Robust Backup Control Barrier Functions: Safety Under Uncertain Dynamics

### v1

@@ IV-B Rigid Body Spacecraft Rotation
Figure 3: Safety-critical control of rigid body spacecraft using the proposed disturbance-robust backup control barrier function approach. (Left) State-space visualization of angular velocity components showing the trajectory of the angular velocity vector over time. (Right) Angular velocity norm over time showing adherence to safety constraint (top), and commanded primary control and actual (safe) control signal over time (bottom).
@@
Consider next an example of a rigid body spacecraft with a known inertia tensor in the body frame given by $\boldsymbol{J}$ where the dynamics of the angular velocities can be described by Euler’s rotational equations of motion
$\boldsymbol{\dot{\omega}}=\boldsymbol{J}^{-1}(-\boldsymbol{\omega}\times\boldsymbol{J}\boldsymbol{\omega}+\boldsymbol{u})+\boldsymbol{d}_{x}.$ (39)
@@
Here, $\boldsymbol{u}$ is the control torque vector that can be applied by the spacecraft to control the angular velocity and $\boldsymbol{d}_{x}$ is an unknown but bounded additive disturbance vector, such that $\left\|\boldsymbol{d}_{x}\right\|\leq\xi\in\mathbb{R}_{>0}$. The states of interest are the angular velocity vector elements in the body frame, $\boldsymbol{\omega}=[\omega_{1},\omega_{2},\omega_{3}]^{T}$.
@@
The safety objective is to ensure that the norm of the angular velocity vector of the spacecraft does not exceed a maximum value. This constraint is designed to prevent damage to sensors that may be onboard the rigid body spacecraft. The safe set is therefore $\mathcal{C}_{\rm S}=\{\boldsymbol{\omega}\in\mathbb{R}^{3}:h(\boldsymbol{\omega})\geq 0\}$ where $h(\boldsymbol{\omega})=\omega_{\rm max}^{2}-\left\|\boldsymbol{\omega}\right\|^{2}\geq 0$ and $\omega_{\rm max}\in\mathbb{R}_{>0}$ represents the maximum allowable angular velocity. The primary controller is given by ${\boldsymbol{u}_{\rm p}(t)=\text{sin}([\frac{t}{2},\frac{t}{2}-\frac{\pi}{4},\frac{t}{4}+\frac{\pi}{4}]^{T})}$. The robust backup control law $\boldsymbol{u}_{\rm b}(\boldsymbol{\omega})=-k_{b}\boldsymbol{J}\boldsymbol{\omega}+\boldsymbol{\omega}\times\boldsymbol{J}\boldsymbol{\omega}$ renders the backup set $\mathcal{C}_{\rm B}\triangleq\{\boldsymbol{\omega}\in\mathbb{R}^{3}:h_{\rm b}(\boldsymbol{\omega})=\gamma-\frac{1}{2}\boldsymbol{\omega}^{T}\boldsymbol{J}\boldsymbol{\omega}\geq 0\}$ robustly forward invariant for sufficiently large gain $k_{b}$. $\mathcal{C}_{\rm B}$ is a level set of the spacecraft’s rotational energy, defined by the scalar $\gamma$. It can be shown that $k_{b}>(\lambda_{\rm max}\xi)/(\sqrt{2\gamma\lambda_{\rm min}})$ ensures $\dot{h}_{\rm b}(\boldsymbol{\omega},\boldsymbol{u})\geq 0$ at the boundary of the level set for (39), where $\lambda_{\rm max}$ and $\lambda_{\rm min}$ are the maximum and minimum eigenvalues of $\boldsymbol{J}$, respectively. The proof is omitted for brevity.
@@
It is straightforward to verify that the closed-loop nominal backup dynamics are strongly contracting with a rate of $k_{b}$ since $f_{\rm cl}(\boldsymbol{\omega})=-k_{b}\boldsymbol{\omega}$. Because the log norm of the closed-loop Jacobian ($\partial f_{\rm cl}/\partial\boldsymbol{\omega}$) is upper bounded by $-k_{b}$, the disturbed and nominal backup flows can be bounded as
@@
by [13, Corollary 3.17]. This yields much tighter $\epsilon_{\tau}$ and $\epsilon_{\rm b}$ terms than the general Gronwall bound.
For the simulations, $\omega_{\rm max}=1\mkern 4.0mu\rm rad/s$, $\gamma=2\rm\mkern 4.0muJ$, $\xi=0.1\mkern 4.0mu{\rm rad/s^{2}}$ and $\boldsymbol{J}=\texttt{diag}\{[12,12,5]\}\mkern 4.0mu{\rm kg\mkern 1.0mum^{2}}$. The $\texttt{diag}\{\cdot\}$ operator returns a diagonal matrix with the inputted vector on the diagonal. The discretization time step used for computing the nominal backup flow was $\Delta=0.05\mkern 4.0mu{\rm s}$ and the integration horizon was $T=1.75\mkern 4.0mu{\rm s}$. The disturbance vector was time-varying, given by $\boldsymbol{d}_{x}(t)=\xi\frac{\boldsymbol{v}(t)}{\|\boldsymbol{v}(t)\|}$ where $\boldsymbol{v}(t)=\text{sin}([\frac{t}{2}+\frac{\pi}{2},\frac{t}{2},\frac{t}{2}-\frac{\pi}{2}]^{T})$.
Figure 3 shows the results of using the disturbance-robust backup CBF method for this application with the contraction bounds in (40). On the left of Figure 3 the trajectory of the disturbed system over time is plotted in blue, and the goal is to keep this trajectory within the red sphere. The magenta sections of the trajectory indicate that the primary control signal, $\boldsymbol{u}_{\rm p}$, has been modified to assure safety. The wire-frame spheres represent the contraction norm balls along the nominal backup flow in cyan. Note that as required by (20), the terminal spheres are entirely contained within the $\mathcal{C}_{\rm B}$ ellipse. The right plots of Figure 3 show that the norm constraint on the angular velocity is adhered to in the presence of time-varying disturbances and that the primary control signal is modified only when $\left\|\boldsymbol{\omega}\right\|$ is near $\omega_{\rm max}$.
@@ V Conclusions
This paper presents a novel safety-critical control framework to handle unknown bounded disturbances for a broad class of nonlinear systems. We extend the method of backup CBFs to handle such disturbances by providing forward invariance conditions for a subset of a controlled invariant set governed by the disturbed system. We prove that enforcing these conditions guarantees safety of the disturbed system and demonstrate the effectiveness of the approach with two numerical simulation examples.

### v2

@@ IV-B Rigid Body Spacecraft Rotation
Figure 3: Simulation results for a rigid body spacecraft, comparing the proposed disturbance-robust backup CBF approach and the standard backup CBF formulation. (Left) State-space visualization of angular velocity components showing the trajectory of the angular velocity vector over time. The objective is to keep the trajectory within the red sphere (safe region). The standard approach violates safety due to the disturbance, while the proposed disturbance-robust method does not. Magenta sections of the blue trajectory indicate that the primary control signal, $\boldsymbol{u}_{\rm p}$, has been modified to assure safety. Wire-frame spheres represent the contraction norm balls along the nominal backup flow in cyan. (Right) Angular velocity norm over time for both approaches (top), and commanded primary control and actual (safe) control signal over time for the robust approach (bottom).
@@
Consider next an example of a rigid body spacecraft with a known inertia tensor in the body frame given by $\boldsymbol{J}$, where the dynamics of the angular velocities can be described by Euler’s rotational equations of motion
$\dot{\boldsymbol{{\omega}}}=\boldsymbol{J}^{-1}(-\boldsymbol{\omega}\times\boldsymbol{J}\boldsymbol{\omega}+\boldsymbol{u})+\boldsymbol{d}_{x}.$ (37)
@@
Here, $\boldsymbol{u}\in\mathcal{U}=[-1,1]^{3}\mkern 4.0mu{\rm Nm}$ is the control torque vector that can be applied by the spacecraft to control the angular velocity and $\boldsymbol{d}_{x}$ is an unknown but bounded additive disturbance vector, such that $\left\|\boldsymbol{d}_{x}\right\|\leq\xi\in\mathbb{R}_{>0}$. The states of interest are the angular velocity vector elements in the body frame.
@@
The safety objective is to ensure that the norm of the angular velocity vector of the spacecraft does not exceed a maximum value, to prevent damage to onboard sensors. The safe set is therefore $\mathcal{C}_{\rm S}=\{\boldsymbol{\omega}\in\mathbb{R}^{3}:h(\boldsymbol{\omega})\geq 0\}$ where $h(\boldsymbol{\omega})=\omega_{\rm max}^{2}-\left\|\boldsymbol{\omega}\right\|^{2}\geq 0$ and $\omega_{\rm max}\in\mathbb{R}_{>0}$ represents the maximum allowable angular velocity. The primary controller is given by ${\boldsymbol{u}_{\rm p}(t)=\text{sin}([\frac{t}{2},\frac{t}{2}-\frac{\pi}{4},\frac{t}{4}+\frac{\pi}{4}]^{T})}$. The robust backup control law $\boldsymbol{u}_{\rm b}(\boldsymbol{\omega})=-k_{b}\boldsymbol{J}\boldsymbol{\omega}+\boldsymbol{\omega}\times\boldsymbol{J}\boldsymbol{\omega}$ renders the backup set $\mathcal{C}_{\rm B}\triangleq\{\boldsymbol{\omega}\in\mathbb{R}^{3}:h_{\rm b}(\boldsymbol{\omega})=\gamma-\frac{1}{2}\boldsymbol{\omega}^{T}\boldsymbol{J}\boldsymbol{\omega}\geq 0\}$ robustly forward invariant for sufficiently large gain $k_{b}$. $\mathcal{C}_{\rm B}$ is a level set of the spacecraft’s rotational energy, defined by the scalar $\gamma$. Any $k_{b}>(\lambda_{\rm max}\xi)/(\sqrt{2\gamma\lambda_{\rm min}})$ ensures $\dot{h}_{\rm b}(\boldsymbol{\omega},\boldsymbol{u})\geq 0$ at the boundary of the level set for (37), where $\lambda_{\rm max}$ and $\lambda_{\rm min}$ are the maximum and minimum eigenvalues of $\boldsymbol{J}$, respectively. The proof is omitted for brevity.
@@
It is straightforward to verify that the closed-loop nominal backup dynamics are strongly contracting with a rate of $k_{b}$ since $f_{\rm cl}(\boldsymbol{\omega})=-k_{b}\boldsymbol{\omega}$. Because the log norm of the closed-loop Jacobian ($\partial f_{\rm cl}/\partial\boldsymbol{\omega}$) is upper-bounded by $-k_{b}$, the disturbed and nominal backup flows can be bounded as
@@
by [28, Corollary 3.17]. This yields tighter $\epsilon_{\tau}$ and $\epsilon_{\rm b}$ terms than the general Gronwall bound.
For the simulations, $\omega_{\rm max}=1\mkern 4.0mu\rm rad/s$, $\gamma=2\rm\mkern 4.0muJ$, $\xi=0.1\mkern 4.0mu{\rm rad/s^{2}}$ and $\boldsymbol{J}$ is diagonal with elements $[12,12,5]\mkern 4.0mu{\rm kg\mkern 1.0mum^{2}}$. The discretization time step used for computing the nominal backup flow is $\Delta=0.05\mkern 4.0mu{\rm s}$ and the integration horizon is $T=1.75\mkern 4.0mu{\rm s}$. The disturbance vector is time-varying, given by $\boldsymbol{d}_{x}(t)=\xi\frac{\boldsymbol{v}(t)}{\|\boldsymbol{v}(t)\|}$ where $\boldsymbol{v}(t)=\text{sin}([\frac{t}{2}+\frac{\pi}{2},\frac{t}{2},\frac{t}{2}-\frac{\pi}{2}]^{T})$.
Figure 3 compares our disturbance-robust backup CBF method using the contraction bounds in (38) with the standard backup CBF approach. Our approach obeys the norm constraint on the angular velocity in the presence of unknown time-varying disturbances, while the standard backup CBF approach does not, violating safety multiple times.
@@ V Conclusions
In this article we presented a novel safety-critical control framework to handle unknown bounded disturbances for a broad class of nonlinear systems. We extended the method of backup CBFs to handle such disturbances by providing forward invariance conditions for a subset of a controlled invariant set governed by the disturbed system. We proved that enforcing these conditions guarantees safety for the disturbed system, and we demonstrated the effectiveness of the approach with two numerical simulation examples.